# Risk spillovers between the BRICS and the U.S. staple grain futures markets

### v1

@@ Abstract
This study examines contemporaneous and lagged spillover effects in BRICS staple grain futures markets and their linkages with U.S. markets. The results show that contemporaneous spillovers dominate, while net spillovers are driven by lagged connectedness. Systemic risk is lower in intra-BRICS markets compared to those including the U.S., highlighting the U.S. grain market’s significant influence. Brazilian and U.S. grains are key net spillover contributors, excluding U.S. rice, while South African staple grains act as major net receivers. Particularly, the spillover between soybeans is the strongest. The study also reveals heterogeneous impacts of the Russia-Ukraine conflict and Black Sea Grain Initiative on grain futures.
@@ 3.1 Average spillover effects
Our analysis starts with the average spillover effects for System 1, as depicted in Fig. 2. The average TCI stands at 41.54%, with 30.55% attributed to contemporaneous effects and 10.99% to lagged dynamics. From the perspectives of overall and contemporaneous connectedness, the spillover of risk between U.S. and Brazilian soybeans is the strongest, aligning with their highest positive correlation, as shown in Fig. 1. Lagged spillover effects are most pronounced between South African soybeans and U.S. and Brazilian soybeans. The TO and FROM indices for U.S. and Brazilian soybeans are the highest, while China’s rice exhibits the lowest. [^1] Values before rounding for TO, FROM, and NET are provided in Table S1 (supplementary results).
@@
The overall TCI for System 2 is 23.56%, as presented in Table S2, with the strongest average spillover between South African soybeans and maize. South African staple grains and Brazilian soybeans contribute the most to spillovers to other markets, while South Africa’s grain market exhibits the strongest FROM spillover. In terms of overall NET connectedness, only Brazilian grains act as net transmitters of shocks. [^2] See supplementary results for detailed intra-BRICS market connectedness.
@@ 3.2 Time-varying total spillover effects
Fig. 3 illustrates the time-varying total connectedness, with the overall and contemporaneous TCI showing similar evolution patterns, and contemporaneous spillover effects markedly exceeding those in the lagged period. The overall TCI of System 1 increased prior to the Russia-Ukraine conflict and began to decline a few days later, while the overall TCI of System 2 slightly dropped immediately after the conflict, followed by minor fluctuations. The TCI of both systems experienced a significant decline after the BSGI extension period was shortened. Clearly, Figs. 2 and 3, along with Table S2, demonstrate that the overall, contemporaneous, and lagged connectedness of System 1 is significantly stronger than that of System 2, highlighting the substantial influence of the U.S. grain futures market on BRICS grain markets.
@@ 3.3 Dynamic directional spillovers: FROM, TO, and NET
Fig. 4 and Fig. S2 depict the time-varying directional connectedness of System 1 and System 2, respectively. At a glance, most grains exhibit similar TO and FROM spillover trajectories, differing mainly in spillover intensity. The outbreak of the Russia-Ukraine conflict and the shortening of the BSGI extension period have inconsistent effects on the TO, FROM, and NET spillover effects across the 18 grain futures. The spillover from China’s and India’s grain futures to other markets is very low, generally below 20%. Particularly, after the outbreak of the conflict, the spillover effects of U.S. and Black Sea wheat on other markets dropped by more than 20%, with the spillover levels remaining stable for an extended period. Additionally, after the BSGI extension was shortened to 60 days, their spillovers also decreased by more than 20%. U.S. corn saw its TO spillover reach 90% for several months after the conflict broke out, before the BSGI came into effect. Compared to U.S. soybeans, corn, and wheat, the TO connectedness of U.S. rice is relatively weak. Therefore, even for different grain futures from the same country, the impact of the Russia-Ukraine conflict varies significantly. Correspondingly, the spillover from other markets received by Indian maize and Chinese rice is also very low, with the overall level remaining below 20%. For System 2, South African and Brazilian grain futures are the primary transmitters and receivers of risk spillovers, followed by Chinese corn. In contrast, Chinese rice has the lowest TO and FROM connectedness throughout the sample period, consistently below 10%.
@@ 3.4 Net pairwise directional spillover effects
From the net pairwise directional spillover of System 1 presented in Fig. 5, it is evident that spillovers between soybean markets are the strongest. Brazil and the U.S. exhibit significant spillover effects to South Africa and China, with the spillover to South Africa reaching as high as 16%. The NPDC between Brazilian and U.S. soybeans changes over time and is significantly influenced by the outbreak of the Russia-Ukraine conflict and the shortening of the BSGI. Additionally, the contemporaneous and lagged NPDC of each grain pair show substantial differences, with the contemporaneous NPDC of BRs and USs being stronger than the lagged one. Overall, the lagged NPDC dominates for most grain pairs. Additionally, most of the NPDC in System 2 follow similar patterns to those in System 1, but with varying spillover intensities. For instance, the NPDC between BRs and ZAs reaches a peak of 16% in System 1, whereas it can reach up to 25% in System 2.333See supplementary results for the remaining NPDC of System 1 and those of System 2.

### v2

@@ Abstract
This study examines contemporaneous and lagged spillovers in BRICS staple grain futures markets and their linkages with U.S. markets. Results show that contemporaneous spillovers dominate, while net spillovers are driven by lagged connectedness. Systemic risk is lower in intra-BRICS markets than in those including the U.S., highlighting the U.S. grain market’s significant influence. Brazilian and U.S. grains are key net spillover contributors, excluding U.S. rice, while South African staple grains act as major net receivers. Notably, spillovers between soybeans are the strongest. The study also reveals heterogeneous impacts of the Russia-Ukraine conflict and Black Sea Grain Initiative on grain futures.
@@ 3.1 Average spillover effects
Our analysis starts with the average spillover effects for System 1, as depicted in Fig. 2. The average TCI stands at 41.54%, with 30.55% attributed to contemporaneous effects and 10.99% to lagged dynamics. From the perspectives of overall and contemporaneous connectedness, the spillover of risk between U.S. and Brazilian soybeans is the strongest, aligning with their highest positive correlation, as shown in Fig. 1. Lagged spillover effects are most pronounced between South African soybeans and U.S. and Brazilian soybeans. The TO and FROM indices for U.S. and Brazilian soybeans are the highest, while China’s rice exhibits the lowest. [^1] Values before rounding for TO, FROM, and NET are provided in Table 2.
@@
The overall TCI for System 2 is 23.56%, as presented in Table 3, with the strongest average spillover between South African soybeans and maize. South African staple grains and Brazilian soybeans contribute the most to spillovers to other markets, while South Africa’s grain market exhibits the strongest FROM spillover. In terms of overall NET connectedness, only Brazilian grains act as net transmitters of shocks.
@@ 3.2 Time-varying total spillover effects
Fig. 3 illustrates the time-varying total connectedness, with the overall and contemporaneous TCI showing similar evolution patterns, and contemporaneous spillover effects markedly exceeding those in the lagged period. The overall TCI of System 1 increased prior to the Russia-Ukraine conflict and began to decline a few days later, while the overall TCI of System 2 slightly dropped immediately after the conflict, followed by minor fluctuations. The TCI of both systems experienced a significant decline after the BSGI extension period was shortened. Clearly, Figs. 2 and 3, along with Table 3, demonstrate that the overall, contemporaneous, and lagged connectedness of System 1 is significantly stronger than that of System 2, highlighting the substantial influence of the U.S. grain futures market on BRICS grain markets.
@@ 3.3 Dynamic directional spillovers: FROM, TO, and NET
Fig. 4 and Fig. 11 depict the time-varying directional connectedness of System 1 and System 2, respectively. At a glance, most grains exhibit similar TO and FROM spillover trajectories, differing mainly in spillover intensity. The outbreak of the Russia-Ukraine conflict and the shortening of the BSGI extension period have inconsistent effects on the TO, FROM, and NET spillover effects across the 18 grain futures. The spillover from China’s and India’s grain futures to other markets is very low, generally below 20%. Particularly, after the outbreak of the conflict, the spillover effects of U.S. and Black Sea wheat on other markets dropped by more than 20%, with the spillover levels remaining stable for an extended period. Additionally, after the BSGI extension was shortened to 60 days, their spillovers also decreased by more than 20%. U.S. corn saw its TO spillover reach 90% for several months after the conflict broke out, before the BSGI came into effect. Compared to U.S. soybeans, corn, and wheat, the TO connectedness of U.S. rice is relatively weak. Therefore, even for different grain futures from the same country, the impact of the Russia-Ukraine conflict varies significantly. Correspondingly, the spillover from other markets received by Indian maize and Chinese rice is also very low, with the overall level remaining below 20%. For System 2, South African and Brazilian grain futures are the primary transmitters and receivers of risk spillovers, followed by Chinese corn. In contrast, Chinese rice has the lowest TO and FROM connectedness throughout the sample period, consistently below 10%.
@@ 3.4 Net pairwise directional spillover effects
From the net pairwise directional spillover of System 1 presented in Fig. 5, it is evident that spillovers between soybean markets are the strongest. Brazil and the U.S. exhibit significant spillover effects to South Africa and China, with the spillover to South Africa reaching as high as 16%. The NPDC between Brazilian and U.S. soybeans changes over time and is significantly influenced by the outbreak of the Russia-Ukraine conflict and the shortening of the BSGI. Additionally, the contemporaneous and lagged NPDC of each grain pair show substantial differences, with the contemporaneous NPDC of BRs and USs being stronger than the lagged one. Overall, the lagged NPDC dominates for most grain pairs. Additionally, most of the NPDC in System 2 follow similar patterns to those in System 1, but with varying spillover intensities. For instance, the NPDC between BRs and ZAs reaches a peak of 16% in System 1, whereas it can reach up to 25% in System 2.
@@ Acknowledgements
Appendix A Appendix
Table 2: Averaged dynamic connectedness of System 1.
BRs CNs ZAs USs BRc INm CNc ZAm BSc USc CNw ZAw BSw USw CNr USr FROM
BRs 0.22 0.79 0.68 39.88 2.44 0.85 1.20 1.00 4.60 5.19 0.65 0.65 2.78 3.66 0.84 1.99 67.19
( 0.00, 0.22) ( 0.34, 0.44) ( 0.57, 0.11) (39.69, 0.19) ( 2.26, 0.18) ( 0.65, 0.20) ( 0.83, 0.37) ( 0.65, 0.35) ( 4.33, 0.26) ( 4.92, 0.27) ( 0.33, 0.32) ( 0.46, 0.19) ( 2.45, 0.33) ( 3.51, 0.16) ( 0.45, 0.39) ( 1.82, 0.17) (63.26, 3.93)
CNs 2.15 0.46 0.93 3.35 1.13 1.10 2.41 0.97 1.13 1.65 1.92 0.83 2.76 0.95 1.62 2.00 24.90
( 0.57, 1.59) ( 0.00, 0.46) ( 0.64, 0.29) ( 0.42, 2.92) ( 0.78, 0.35) ( 0.54, 0.56) ( 2.10, 0.31) ( 0.74, 0.24) ( 0.25, 0.88) ( 0.99, 0.66) ( 1.78, 0.14) ( 0.46, 0.37) ( 0.29, 2.47) ( 0.37, 0.57) ( 0.42, 1.20) ( 0.21, 1.79) (10.57, 14.33)
ZAs 9.56 0.95 1.51 10.34 2.46 0.69 2.88 18.10 1.72 2.06 0.43 6.47 1.63 1.36 0.59 1.45 60.70
( 0.56, 9.00) ( 0.57, 0.38) ( 0.00, 1.51) ( 0.79, 9.55) ( 0.46, 1.99) ( 0.41, 0.28) ( 2.15, 0.73) (17.42, 0.68) ( 0.55, 1.16) ( 0.25, 1.81) ( 0.26, 0.18) ( 6.09, 0.38) ( 0.80, 0.83) ( 0.52, 0.84) ( 0.23, 0.36) ( 0.41, 1.04) (31.47, 29.23)
USs 39.50 0.85 0.91 0.36 1.79 0.97 1.15 0.87 4.34 6.57 2.14 0.62 2.56 3.45 0.69 2.09 68.50
(39.09, 0.41) ( 0.25, 0.59) ( 0.82, 0.10) ( 0.00, 0.36) ( 1.61, 0.18) ( 0.60, 0.37) ( 0.84, 0.32) ( 0.62, 0.25) ( 4.01, 0.34) ( 6.23, 0.34) ( 0.58, 1.56) ( 0.39, 0.23) ( 2.01, 0.55) ( 3.09, 0.36) ( 0.56, 0.13) ( 1.58, 0.51) (62.26, 6.24)
BRc 2.89 0.99 1.21 1.94 0.97 1.65 0.83 2.18 3.36 4.48 0.67 1.47 1.05 2.50 0.64 1.56 27.43
( 2.45, 0.45) ( 0.74, 0.24) ( 0.54, 0.67) ( 1.62, 0.32) ( 0.00, 0.97) ( 0.90, 0.75) ( 0.53, 0.30) ( 1.65, 0.53) ( 2.88, 0.48) ( 3.66, 0.82) ( 0.16, 0.51) ( 0.47, 1.01) ( 0.78, 0.27) ( 2.22, 0.29) ( 0.41, 0.23) ( 1.00, 0.57) (20.00, 7.43)
INm 1.04 0.83 1.29 1.03 1.64 2.02 1.66 1.64 0.87 0.64 0.86 0.69 1.00 0.70 0.76 1.22 15.87
( 0.77, 0.27) ( 0.56, 0.28) ( 0.59, 0.70) ( 0.69, 0.34) ( 0.97, 0.67) ( 0.00, 2.02) ( 0.89, 0.77) ( 0.86, 0.78) ( 0.17, 0.69) ( 0.23, 0.41) ( 0.24, 0.63) ( 0.22, 0.47) ( 0.36, 0.64) ( 0.32, 0.38) ( 0.25, 0.51) ( 0.49, 0.73) ( 7.62, 8.26)
CNc 1.76 2.24 3.04 1.53 1.56 1.25 0.45 2.37 2.02 3.77 0.83 1.38 1.94 1.48 1.63 0.99 27.78
( 0.92, 0.84) ( 2.07, 0.18) ( 2.46, 0.58) ( 0.91, 0.61) ( 0.52, 1.04) ( 0.84, 0.41) ( 0.00, 0.45) ( 1.59, 0.78) ( 1.43, 0.60) ( 0.38, 3.39) ( 0.67, 0.15) ( 0.93, 0.44) ( 1.20, 0.74) ( 0.71, 0.77) ( 1.10, 0.53) ( 0.30, 0.70) (16.03, 11.76)
ZAm 2.59 0.91 18.59 2.79 3.62 0.92 1.95 0.34 3.09 5.12 0.79 10.53 1.77 2.17 0.34 1.03 56.20
( 0.76, 1.83) ( 0.62, 0.29) (18.20, 0.39) ( 0.72, 2.07) ( 1.45, 2.17) ( 0.61, 0.31) ( 1.53, 0.42) ( 0.00, 0.34) ( 1.15, 1.94) ( 0.41, 4.70) ( 0.31, 0.48) (10.17, 0.36) ( 0.56, 1.21) ( 0.75, 1.42) ( 0.15, 0.19) ( 0.51, 0.51) (37.90, 18.30)
BSc 4.68 0.77 1.77 4.24 3.02 0.65 1.55 1.62 3.80 8.84 0.91 0.70 12.02 10.70 0.50 2.54 54.51
( 4.17, 0.51) ( 0.16, 0.61) ( 0.54, 1.24) ( 3.74, 0.50) ( 2.60, 0.42) ( 0.10, 0.55) ( 1.25, 0.30) ( 1.14, 0.48) ( 0.00, 3.80) ( 8.22, 0.62) ( 0.66, 0.25) ( 0.38, 0.32) (11.58, 0.44) (10.35, 0.35) ( 0.18, 0.32) ( 1.99, 0.55) (47.06, 7.45)
USc 5.19 1.28 0.62 6.77 3.67 1.44 0.66 0.92 8.95 1.65 1.15 1.12 5.49 6.32 0.56 2.05 46.19
( 4.68, 0.50) ( 0.84, 0.45) ( 0.31, 0.31) ( 6.16, 0.60) ( 3.31, 0.36) ( 0.17, 1.28) ( 0.33, 0.34) ( 0.46, 0.46) ( 8.36, 0.58) ( 0.00, 1.65) ( 0.76, 0.40) ( 0.70, 0.41) ( 4.80, 0.68) ( 5.24, 1.08) ( 0.29, 0.27) ( 1.74, 0.31) (38.16, 8.03)
CNw 0.81 2.01 0.75 1.10 0.66 0.94 1.07 1.09 1.30 1.22 0.56 2.74 1.30 2.95 0.66 0.97 19.57
( 0.44, 0.38) ( 1.85, 0.16) ( 0.36, 0.39) ( 0.80, 0.30) ( 0.18, 0.48) ( 0.24, 0.70) ( 0.69, 0.38) ( 0.38, 0.70) ( 0.87, 0.43) ( 0.91, 0.31) ( 0.00, 0.56) ( 1.79, 0.95) ( 0.58, 0.73) ( 0.89, 2.06) ( 0.45, 0.21) ( 0.38, 0.59) (10.80, 8.77)
ZAw 1.45 0.53 6.85 1.55 1.18 0.29 1.14 10.67 3.64 2.73 2.20 2.55 5.17 7.07 0.67 1.69 46.82
( 0.58, 0.87) ( 0.32, 0.20) ( 6.55, 0.29) ( 0.52, 1.03) ( 0.42, 0.76) ( 0.14, 0.16) ( 0.87, 0.27) (10.25, 0.42) ( 0.41, 3.23) ( 0.70, 2.03) ( 1.53, 0.67) ( 0.00, 2.55) ( 2.36, 2.82) ( 1.18, 5.89) ( 0.37, 0.29) ( 1.08, 0.61) (27.28, 19.54)
BSw 3.37 0.81 1.50 2.46 0.94 0.57 1.16 1.42 13.33 5.36 1.10 2.88 0.79 16.36 1.12 1.09 53.46
( 2.56, 0.82) ( 0.22, 0.59) ( 0.87, 0.62) ( 2.03, 0.43) ( 0.69, 0.25) ( 0.27, 0.31) ( 0.97, 0.19) ( 0.63, 0.79) (11.89, 1.43) ( 4.96, 0.40) ( 0.46, 0.64) ( 2.48, 0.39) ( 0.00, 0.79) (15.69, 0.67) ( 0.56, 0.57) ( 0.77, 0.32) (45.04, 8.43)
USw 3.95 0.48 0.84 3.22 2.24 0.43 1.05 1.19 11.28 6.24 0.98 2.46 16.13 0.49 0.55 3.40 54.42
( 3.48, 0.47) ( 0.29, 0.19) ( 0.54, 0.30) ( 2.90, 0.31) ( 2.01, 0.23) ( 0.21, 0.22) ( 0.62, 0.42) ( 0.79, 0.40) (10.60, 0.68) ( 5.19, 1.05) ( 0.73, 0.25) ( 1.12, 1.34) (15.59, 0.54) ( 0.00, 0.49) ( 0.21, 0.34) ( 2.82, 0.58) (47.10, 7.32)
CNr 0.96 1.13 1.10 1.21 0.99 0.83 1.60 0.56 0.69 0.57 0.84 1.07 1.33 1.00 4.36 0.68 14.55
( 0.56, 0.40) ( 0.44, 0.69) ( 0.36, 0.74) ( 0.71, 0.50) ( 0.44, 0.55) ( 0.24, 0.58) ( 1.15, 0.45) ( 0.23, 0.33) ( 0.25, 0.44) ( 0.38, 0.19) ( 0.45, 0.39) ( 0.45, 0.63) ( 0.71, 0.62) ( 0.30, 0.71) ( 0.00, 4.36) ( 0.39, 0.29) ( 7.04, 7.51)
USr 2.86 0.88 1.20 2.87 1.72 1.46 0.51 0.96 2.93 2.58 1.15 1.69 1.36 3.67 0.72 1.74 26.56
( 2.07, 0.78) ( 0.21, 0.67) ( 0.54, 0.67) ( 1.78, 1.09) ( 1.01, 0.71) ( 0.45, 1.02) ( 0.29, 0.22) ( 0.60, 0.36) ( 2.20, 0.73) ( 1.96, 0.61) ( 0.36, 0.79) ( 1.32, 0.37) ( 0.83, 0.53) ( 3.18, 0.49) ( 0.37, 0.35) ( 0.00, 1.74) (17.18, 9.38)
TO 82.76 15.44 41.27 84.27 29.06 14.05 20.82 45.55 63.24 57.01 16.62 35.29 58.29 64.36 11.87 24.75 664.66
( 63.66, 19.11) ( 9.47, 5.97) ( 33.87, 7.40) ( 63.49, 20.77) ( 18.73, 10.33) ( 6.35, 7.70) ( 15.04, 5.78) ( 38.00, 7.55) ( 49.35, 13.89) ( 39.41, 17.61) ( 9.27, 7.35) ( 27.42, 7.87) ( 44.89, 13.40) ( 48.32, 16.04) ( 6.00, 5.88) ( 15.49, 9.27) (488.75, 175.91)
Inc.Own 82.98 15.90 42.78 84.62 30.03 16.07 21.27 45.89 67.04 58.67 17.18 37.84 59.08 64.85 16.23 26.50 TCI
(63.66, 19.32) ( 9.47, 6.43) (33.87, 8.91) (63.49, 21.13) (18.73, 11.30) ( 6.35, 9.72) (15.04, 6.23) (38.00, 7.89) (49.35, 17.69) (39.41, 19.26) ( 9.27, 7.91) (27.42, 10.42) (44.89, 14.19) (48.32, 16.53) ( 6.00, 10.23) (15.49, 11.01) ($\rm TCI^{C}$, $\rm TCI^{L}$)
NET 15.57 -9.46 -19.43 15.77 1.63 -1.83 -6.96 -10.65 8.73 10.82 -2.95 -11.53 4.83 9.94 -2.68 -1.81 41.54
( 0.39, 15.18) (-1.10, -8.36) ( 2.39, -21.82) ( 1.23, 14.54) (-1.27, 2.90) (-1.27, -0.56) (-0.99, -5.97) ( 0.10, -10.75) ( 2.29, 6.44) ( 1.25, 9.57) (-1.53, -1.42) ( 0.14, -11.67) (-0.15, 4.98) ( 1.22, 8.72) (-1.04, -1.64) (-1.69, -0.12) (30.55, 10.99)
Table 3: Averaged dynamic connectedness of intra-BRICS markets.
BRs CNs ZAs BRc INm CNc ZAm CNw ZAw CNr FROM
BRs 0.46 0.98 1.32 6.71 1.47 2.39 1.46 1.21 0.85 1.43 17.83
( 0.00, 0.46) ( 0.54, 0.44) ( 1.14, 0.18) ( 6.41, 0.30) ( 1.20, 0.27) ( 1.69, 0.70) ( 0.82, 0.64) ( 0.62, 0.59) ( 0.45, 0.40) ( 0.83, 0.60) (13.72, 4.11)
CNs 3.78 0.47 1.19 1.06 1.03 2.43 1.19 1.84 0.46 1.60 14.58
( 0.56, 3.22) ( 0.00, 0.47) ( 0.96, 0.23) ( 0.72, 0.34) ( 0.59, 0.44) ( 2.07, 0.35) ( 0.99, 0.20) ( 1.69, 0.15) ( 0.26, 0.20) ( 0.58, 1.02) ( 8.42, 6.16)
ZAs 16.15 1.25 1.78 2.72 0.77 3.31 21.08 0.42 7.11 0.60 53.42
( 0.97, 15.19) ( 0.88, 0.38) ( 0.00, 1.78) ( 0.49, 2.23) ( 0.44, 0.33) ( 2.42, 0.90) (20.46, 0.62) ( 0.28, 0.14) ( 6.71, 0.40) ( 0.21, 0.39) (32.85, 20.57)
BRc 6.85 0.97 1.32 0.89 1.75 1.11 2.41 0.68 1.71 0.74 17.53
( 6.44, 0.40) ( 0.70, 0.27) ( 0.55, 0.78) ( 0.00, 0.89) ( 0.90, 0.85) ( 0.78, 0.32) ( 1.93, 0.47) ( 0.17, 0.51) ( 0.69, 1.02) ( 0.46, 0.28) (12.62, 4.90)
INm 1.54 0.88 1.33 1.58 2.13 1.64 1.64 0.81 0.69 0.77 10.88
( 1.22, 0.32) ( 0.60, 0.28) ( 0.59, 0.75) ( 0.94, 0.65) ( 0.00, 2.13) ( 0.89, 0.75) ( 0.78, 0.85) ( 0.23, 0.58) ( 0.19, 0.50) ( 0.22, 0.55) ( 5.65, 5.23)
CNc 2.84 2.21 3.37 2.19 1.30 0.49 2.95 0.79 1.65 1.58 18.88
( 1.66, 1.17) ( 2.06, 0.15) ( 2.77, 0.60) ( 0.78, 1.41) ( 0.86, 0.44) ( 0.00, 0.49) ( 2.19, 0.77) ( 0.68, 0.11) ( 1.28, 0.38) ( 1.04, 0.53) (13.32, 5.56)
ZAm 4.11 1.12 21.59 4.63 0.89 2.47 0.28 0.76 14.27 0.33 50.17
( 0.70, 3.41) ( 0.86, 0.26) (21.14, 0.45) ( 1.69, 2.94) ( 0.59, 0.30) ( 2.03, 0.44) ( 0.00, 0.28) ( 0.28, 0.48) (13.74, 0.53) ( 0.14, 0.19) (41.18, 8.99)
CNw 0.93 1.88 0.81 0.78 1.00 1.08 1.05 0.47 3.19 0.66 11.40
( 0.64, 0.29) ( 1.72, 0.16) ( 0.37, 0.43) ( 0.17, 0.61) ( 0.23, 0.77) ( 0.70, 0.39) ( 0.33, 0.72) ( 0.00, 0.47) ( 2.26, 0.94) ( 0.48, 0.18) ( 6.90, 4.50)
ZAw 1.79 0.42 7.58 1.56 0.34 1.70 15.04 2.48 3.42 0.74 31.64
( 0.37, 1.41) ( 0.22, 0.20) ( 7.25, 0.34) ( 0.64, 0.92) ( 0.14, 0.21) ( 1.25, 0.45) (14.61, 0.43) ( 1.99, 0.48) ( 0.00, 3.42) ( 0.46, 0.28) (26.93, 4.72)
CNr 1.10 1.23 0.96 1.14 0.75 1.51 0.51 0.85 1.23 4.26 9.27
( 0.84, 0.25) ( 0.59, 0.64) ( 0.29, 0.67) ( 0.47, 0.66) ( 0.22, 0.54) ( 1.08, 0.43) ( 0.20, 0.31) ( 0.47, 0.37) ( 0.52, 0.71) ( 0.00, 4.26) ( 4.69, 4.59)
TO 39.09 10.94 39.47 22.37 9.32 17.65 47.33 9.84 31.15 8.45 235.61
( 13.42, 25.67) ( 8.17, 2.78) ( 35.06, 4.41) ( 12.30, 10.06) ( 5.17, 4.15) ( 12.91, 4.74) ( 42.31, 5.02) ( 6.43, 3.41) ( 26.09, 5.06) ( 4.42, 4.03) (166.28, 69.33)
Inc.Own 39.55 11.41 41.25 23.25 11.45 18.14 47.60 10.30 34.57 12.71 TCI
(13.42, 26.13) ( 8.17, 3.25) (35.06, 6.19) (12.30, 10.95) ( 5.17, 6.28) (12.91, 5.23) (42.31, 5.29) ( 6.43, 3.88) (26.09, 8.48) ( 4.42, 8.29) ($\rm TCI^{C}$, $\rm TCI^{L}$)
NET 21.26 -3.64 -13.95 4.84 -1.56 -1.23 -2.85 -1.56 -0.49 -0.82 23.56
(-0.30, 21.56) (-0.25, -3.38) ( 2.21, -16.16) (-0.32, 5.16) (-0.48, -1.08) (-0.41, -0.82) ( 1.12, -3.97) (-0.47, -1.09) (-0.84, 0.34) (-0.26, -0.56) (16.63, 6.93)
  • Notes: $R^{2}$ decomposed measures are based on a 200-day rolling-window VAR model with a lag length of order one (BIC). Values in parentheses represent contemporaneous and lagged effects, respectively.
Figure 8: Continue…
Figure 9: Continue…
Figure 10: Net pairwise directional connectedness of System 1.
Figure 11: Directional connectedness. Panels (a), (b), and (c) represent the TO, FROM, and NET connectedness of intra-BRICS markets (System 2), respectively.
Figure 12: Net pairwise directional connectedness of System 2.
Figure 13: Net pairwise directional spillover networks of System 2. Notes: Edge thickness represents spillover strength. Blue and yellow nodes are transmitters and receivers, respectively. Panels (a) to (d) correspond to the full sample, Subsample 1, Subsample 2, and Subsample 3, showing overall, contemporaneous, and lagged dependency from left to right.